### v1

@@ Abstract
Climate change poses critical challenges globally, disproportionately affecting low-income countries that often lack resources and linguistic representation on the international stage. Despite Bangladesh’s status as one of the most vulnerable nations to climate impacts, research gaps persist in Bengali-language studies related to climate change and NLP. To address this disparity, we introduce dhoroNii (Dhoroni, a novel Bengali (Bangla) climate change and environmental news dataset, comprising a 2300 annotated Bangla news articles, offering multiple perspectives such as political influence, scientific/statistical data, authenticity, stance detection, and stakeholder involvement. Furthermore, we present an in-depth exploratory analysis of Dhoroni and introduce BanglaBERT-Dhoroni family, a novel baseline family for climate stance detection in Bangla, fine-tuned on our dataset. This research contributes significantly to enhancing accessibility and analysis of climate discourse in Bengali (Bangla), addressing crucial communication and research gaps in climate-impacted regions like Bangladesh with 180 million people.
keywords:
[…]
{highlights}
We introduce Dhoroni, a novel benchmark dataset with 2,300 Bengali news articles.
The dataset is annotated by three annotators across ten different perspectives.
Detailed exploratory analysis and reasoning for each perspective are provided.
We present ten baseline models under BanglaBERT-Dhoroni family for different tasks.
Benchmarking scores show stable performance in different tasks.
@@ 1 Introduction
Figure 1: Multiple perspectives annotated and discussed in dhoroNii (Dhoroni) dataset.
The spread of news and articles is crucial for shaping public understanding and awareness of climate change and environmental pollution. Media coverage plays a vital role in informing the public, driving policy decisions, and mobilizing climate action Hase et al. (2021); Olausson (2009); Nassanga et al. (2016). However, this vital communication process faces significant challenges due to the widespread dissemination of misinformation and disinformation. These issues are exacerbated by the overwhelming amount of content generated online, leading to information overload. In this saturated environment, manipulative and misleading content often circulates unchecked, creating confusion in public debates and distorting the science behind climate change (Hase et al., 2021). This confusion contributes to political inaction, as conflicting narratives slow down decision-making processes, dilute the urgency of the issue, and hinder the implementation of effective climate policies (Benegal and Scruggs, 2018; Ding et al., 2011). Misinformation, whether intentional or accidental, has the power to shape public perceptions, leading to skepticism or denial of climate change, ultimately weakening collective action efforts (Olausson, 2009).
In this context, Natural Language Processing (NLP) emerges as a critical tool for addressing these challenges. By enabling the automated processing, analysis, and classification of vast amounts of textual data, NLP can help identify, filter, and correct misleading information while amplifying scientifically accurate content. This becomes especially important in low-resource language regions, such as Bengali (Bangla)-speaking communities, where access to accurate climate information is limited. Through NLP, improved information dissemination can reach the populations most affected by climate change, ensuring that they have the knowledge necessary to advocate for and implement more effective climate policies (Stede and Patz, 2021; Nicholls and Culpepper, 2020). NLP can also enhance the accessibility of global climate discussions, helping to break down language barriers that prevent these communities from engaging in critical debates and receiving the latest scientific developments (Nicholls and Culpepper, 2020).
@@
To overcome these issues, we present dhoroNii, a novel multi perspective bangla climate change and environmental news dataset. dhoroNii (pronounced "Dhara n . ī" and written as "Dhoroni" in English alphabet) signifies "mother earth" in Bengali (Bangla), inspiring us to preserve our birthplace, Mother Earth, and protect the environment to prevent climate change. In addition to the dataset, we introduce a fmaily of baseline Bengali (Bangla) language models, named BanglaBERT-Dhoroni, fine-tuned on this dataset. This research gap creates a significant barrier to the effective processing and sharing of climate-related information from Bangladesh within the global scientific community. By underutilizing the potential of NLP for languages like Bengali (Bangla), the international conversation surrounding climate change risks becoming skewed, favoring voices from more developed, English-speaking nations. This further delays the recognition and action needed to address the environmental crises faced by nations like Bangladesh. Therefore, it is imperative to prioritize developing NLP technologies for low-resource languages to ensure a more equitable and inclusive approach to global climate solutions.
@@
    We introduce dhoroNii (Dhoroni), a novel multi-perspective benchmark dataset, containing a collection of 2300 Bengali (Bangla) news articles from various sources and perspectives worldwide.
@@ 2 Related Works
The intersection of climate change and NLP has become increasingly important, with researchers exploring how NLP can contribute to understanding and addressing climate challenges. Stede and Patz (2021); Rolnick et al. (2022) emphasize the significance of integrating NLP with social science to navigate complex climate discussions, providing valuable insights for policy-making and activism. Mallick et al. (2024) highlight the role of NLP in understanding regional climate impacts, stressing the importance of localized adaptation and mitigation efforts. Furthermore, studies on climate awareness in NLP research emphasize the importance of fostering climate-conscious approaches within the NLP community, advocating for technology-driven solutions that contribute to global climate action (Hershcovich et al., 2022).
@@
Figure 2: Keyword used for data collection in dhoroNii (Dhoroni) dataset.
@@ 4 Model Design and Development
Model.  We have fine-tuned BanglaBERT (sagorsarker/bangla-bert-base [^2] https://huggingface.co/sagorsarker/bangla-bert-base [^3] https://github.com/sagorbrur/bangla-bert, which is freely available in Hugging Face by transformers library (Wolf et al., 2020) with MIT Licence. It is a pretrained language model of Bengali language using mask language modeling described in BERT (Devlin et al., 2019).
@@
Task Model Name Location
Stance Detection BanglaBERT-Dhoroni-SD ciol-research/BanglaBERT-Dhoroni-SD
News Authenticity Identification BanglaBERT-Dhoroni-NAD ciol-research/BanglaBERT-Dhoroni-NAD
Political Influence Detection BanglaBERT-Dhoroni-PAD ciol-research/BanglaBERT-Dhoroni-PAD
Scientific Data Usage Detection BanglaBERT-Dhoroni-SDU ciol-research/BanglaBERT-Dhoroni-SDU
Statistical Data Usage Detection BanglaBERT-Dhoroni-StDU ciol-research/BanglaBERT-Dhoroni-StDU
Data Sources Detection BanglaBERT-Dhoroni-DSD ciol-research/BanglaBERT-Dhoroni-DSD
Impacted Location or Countries Detection BanglaBERT-Dhoroni-ILD ciol-research/BanglaBERT-Dhoroni-ILD
Detecting Climate/Environmental Topics BanglaBERT-Dhoroni-CET ciol-research/BanglaBERT-Dhoroni-CET
News Target Detection BanglaBERT-Dhoroni-NTD ciol-research/BanglaBERT-Dhoroni-NTD
Authority Involvement Detection BanglaBERT-Dhoroni-AID ciol-research/BanglaBERT-Dhoroni-AID
@@ 5.1 Experimental Results and Discussion
No Task Accuracy ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$) F1 Score ($\uparrow$)
@@
7 Impacted Location or Countries Detection 0.536 0.138 0.175 0.134
8 Detecting Climate/Environmental Topics 0.052 0.015 0.076 0.017
@@
In Stance Detection, the accuracy of 0.635 suggests the model performs reasonably well overall, but the low precision (0.426) and recall (0.295) indicate it struggles to correctly identify stances, especially minority or less obvious ones. Stances, such as "support" or "oppose," can be subtle in language, and the model likely finds it difficult to differentiate between them, leading to more incorrect classifications. Moving on to News Authenticity Identification, while the accuracy is impressively high at 0.911, precision (0.184) and recall (0.198) are extremely low. This is a classic sign of class imbalance: the model may be biased toward classifying most news as authentic, since it’s likely the majority class in the training data. As a result, it rarely identifies false news, which heavily skews its precision and recall. A more balanced dataset or targeted strategies to boost minority class detection could help improve these scores. In Political Influence Detection, we see a better balance between precision (0.417) and recall (0.500), which is reflected in a decent F1 score of 0.455. This suggests the model is able to identify political influence with moderate success, though it still tends to miss some instances or incorrectly classify others. Political influence might vary significantly depending on context, which could explain why the model sometimes struggles to capture it consistently. Scientific Data Usage Detection performs relatively well, with an accuracy of 0.870 and a balanced F1 score of 0.465. The model seems to manage identifying scientific data usage, likely because this task involves more structured or identifiable cues (like specific terminology or phrasing). However, the model’s slight tendency to overestimate scientific usage (higher recall than precision) may lead it to incorrectly flag some cases. In contrast, Statistical Data Usage Detection has lower accuracy (0.594), though its precision (0.491) and recall (0.497) are more balanced. This suggests the model can identify statistical data usage somewhat effectively, but the overall complexity of the task—where statistical data may appear in various forms—leads to less consistent results. The slight drop in F1 score (0.440) points to misclassification of some borderline cases. Data Sources Detection shows moderate performance with an accuracy of 0.623, and relatively balanced precision and recall (both around 0.48). This task may be easier for the model to manage because references to data sources are often explicit, making it easier to pick up on relevant patterns. However, there may still be subtle variations in how sources are mentioned, leading to some misclassifications. On the other hand, Impacted Location or Countries Detection shows poor performance. With a precision of 0.138 and an F1 score of 0.134, it’s clear the model struggles to consistently identify locations or countries affected by events. This might be due to the broad range of ways in which locations are mentioned in text, from indirect references to unclear geographical descriptions, making it harder for the model to extract this information reliably. A particularly striking failure is seen in Detecting Climate/Environmental Topics, where the accuracy is just 0.052, and both precision and recall are near zero. This suggests the model is almost completely unable to identify climate-related content, likely due to either insufficient training data or highly ambiguous definitions of what constitutes a "climate" topic. It could also be that the model is not tuned to capture the wide variety of terms and context in which climate issues are discussed. News Target Detection also shows significant challenges, with low precision (0.048) and a recall of only 0.125. Identifying the target of news content is inherently complex, often involving subtle contextual cues that the model seems unable to consistently capture. This might explain the low F1 score of 0.069, as the model struggles to pinpoint the correct targets amidst varying contexts. Lastly, Authority Involvement Detection presents similarly poor results, with an accuracy of 0.183 and extremely low precision (0.030). Authority involvement may be indicated in a variety of ways—sometimes indirectly—which the model likely fails to grasp. The low recall (0.102) further confirms that the model is not identifying enough relevant instances, perhaps due to unclear definitions of authority involvement in the data.
@@ 6 Discussion
As awareness of environmental issues increases among governments and the public (Liu et al., 2023), our work with the Dhoroni dataset and the BanglaBERT-Dhoroni model family contributes significantly to this vital discourse. By providing a comprehensive resource that captures climate-related narratives in Bengali, we aim to facilitate informed discussions on sustainable environmental management. Our efforts not only empower environmental managers and stakeholers, but also engage a wider audience, enabling them to better understand and respond to the complexities of climate change (Ghimire and Panday, 2016) and its local implications.
@@ 6.1 Societal Impact
Societal impact of the Dhoroni dataset cannot be overstated. In Bangladesh, where climate change poses significant threats (Nishat and Mukherjee, 2013), understanding public discourse around these issues is crucial for driving social change (Bell, 1994). The dataset offers a platform for analyzing how climate-related information is communicated and perceived within the community. This understanding can empower citizens to advocate for more robust climate policies and hold leaders accountable for environmental actions.
By providing a rich repository of news articles, the dataset encourages grassroots movements focused on climate awareness and action. Local organizations can harness insights derived from the dataset to mobilize communities, fostering a collective sense of responsibility toward environmental conservation. This grassroots approach can create a ripple effect, leading to increased public engagement and advocacy for climate-related policies. Furthermore, the dataset’s emphasis on authenticity and data sources plays a pivotal role in combating misinformation. In an age where misinformation can easily spread through social media and news outlets (Allcott et al., 2019), having a reliable dataset to verify claims becomes essential. By highlighting credible sources and assessing the authenticity of information, the dataset can serve as a foundation for educating the public about the importance of fact-checking and responsible consumption of news (López-Marcos and Vicente-Fernández, 2021).
@@
Societal implications of the dataset also extend to gender and social equity issues. Climate change disproportionately affects marginalized communities, including women (Islam and Winkel, 2017), low-income households (Haque et al., 2014), and indigenous populations (Abate and Kronk, 2013). By analyzing media representations of these groups in the context of climate change, researchers can identify gaps in coverage and advocate for more inclusive narratives. This focus on equity is essential for ensuring that all voices are heard in the conversation about climate change and its impacts. Ultimately, the Dhoroni dataset has the potential to catalyze positive societal change by fostering informed discussions, encouraging activism, and promoting equity in climate discourse. By leveraging the insights generated from this dataset, individuals and organizations can work together to address climate challenges and create a more sustainable future.
@@ 6.2 Global Impact
Global impact of the Dhoroni dataset extends beyond the borders of Bangladesh, contributing to the broader conversation on climate change and environmental issues. As climate change is a global challenge (Bhat and Lone, 2023), understanding how different countries and cultures discuss these topics is vital. By providing a dataset focused on Bengali perspectives, the Dhoroni project adds an important dimension to the global discourse, highlighting the voices and experiences of a population that is often overlooked in international discussions.
The inclusion of Bengali-language content in global climate research is crucial for creating a more balanced narrative. Much of the existing climate-related research and discourse is dominated by English-language sources (Ghosh and Boykoff, 2019), which can skew the understanding of how climate change affects various regions and communities (Leal Filho et al., 2022). By amplifying Bengali voices, the dataset helps to ensure that the unique challenges and perspectives of Bangladeshis are recognized in global climate discussions. Additionally, the dataset can facilitate cross-cultural comparisons in climate reporting and public perception. Researchers and policymakers can analyze how climate change is framed in Bangladesh compared to other countries, revealing potential differences in priorities and narratives. This comparative analysis can inform international collaboration on climate initiatives, allowing for a more holistic understanding of the global climate landscape. On a global scale, BanglaBERT-Dhoroni family models can enhance understanding of how different regions are affected by news topics, such as environmental impacts or political issues. By detecting authenticity and target groups, they can help gauge the global relevance of news stories (Gurevitch and Kavoori, 1994) and ensure that critical issues, such as authority involvement and environmental topics, are accurately represented across different parts of the world.
@@
6.3 Dhoroni and Environmental Management
Our Dhoroni dataset represents a crucial advancement in the intersection of natural language processing and environmental management, particularly for Bengali-speaking communities. By compiling 2,300 news articles that cover diverse perspectives on climate change, the dataset offers a rich resource for understanding how environmental issues are framed in local media. This understanding is vital for effective environmental management, as public perception and discourse significantly influence policy-making and community engagement in climate action (Sofiullahi et al., 2024). BanglaBERT-Dhoroni family model family, fine-tuned on this dataset, enables proper analysis and understanding of climate-related content. By leveraging the powerful capabilities of transformer-based architectures, these models can discern subtle patterns and relationships within the text (Khan et al., 2022), facilitating better detection of climate topics and authority involvement in environmental issues. Such insights are essential for stakeholders in environmental management, including policymakers, researchers, and NGOs, as they seek to tailor their strategies to the unique context of Bengali-speaking regions.
Moreover, the use of advanced models allows for more effective monitoring of public sentiment and discourse regarding climate change (Chen et al., 2019). By analyzing how climate issues are discussed in the media, stakeholders can identify areas where public understanding is lacking and develop targeted communication strategies to address misinformation or raise awareness. Additionally, the ability to classify articles based on various perspectives aids in the identification of gaps in coverage or representation, ensuring that marginalized voices are heard in climate discourse. Overall, we believe Dhoroni dataset and its associated model family can empower researchers and practitioners in environmental management to foster a more informed and engaged public. By enhancing the accessibility and understanding of climate issues through NLP, this initiative contributes to building a more resilient and proactive community in addressing the pressing challenges posed by climate change.

### v2

@@ Abstract
Climate change poses critical challenges globally, disproportionately affecting low-income countries that often lack resources and linguistic representation on the international stage. Despite Bangladesh’s status as one of the most vulnerable nations to climate impacts, research gaps persist in Bengali-language studies related to climate change and NLP. To address this disparity, we introduce Dhoroni, a novel Bengali (Bangla) climate change and environmental news dataset, comprising a 2300 annotated Bangla news articles, offering multiple perspectives such as political influence, scientific / statistical data, authenticity, position detection, and stakeholder involvement. Furthermore, we present an in-depth exploratory analysis of Dhoroni and introduce BanglaBERT-Dhoroni family, a novel baseline model family for climate and environmental opinion detection in Bangla, fine-tuned on our dataset. This research contributes significantly to enhancing accessibility and analysis of climate discourse in Bengali (Bangla), addressing crucial communication and research gaps in climate-impacted regions like Bangladesh with 180 million people.
@@ 1 Introduction
Figure 1: Multiple perspectives annotated and discussed in Dhoroni dataset.
The spread of news and articles is crucial for shaping public understanding and awareness of climate change and environmental pollution. Media coverage plays a vital role in informing the public, driving policy decisions, and mobilizing climate action Hase et al. (2021); Olausson (2009); Nassanga et al. (2016). However, this vital communication process faces significant challenges due to the widespread dissemination of misinformation and disinformation. These issues are exacerbated by the overwhelming amount of content generated online, leading to information overload. In this saturated environment, manipulative and misleading content often circulates unchecked, creating confusion in public debates and distorting the science behind climate change (Hase et al., 2021). This confusion contributes to political inaction, as conflicting narratives slow down decision-making processes, dilute the urgency of the issue, and hinder the implementation of effective climate policies (Benegal & Scruggs, 2018; Ding et al., 2011). Misinformation, whether intentional or accidental, has the power to shape public perceptions, leading to skepticism or denial of climate change, ultimately weakening collective action efforts (Olausson, 2009).
In this context, Natural Language Processing (NLP) emerges as a critical tool for addressing these challenges. By enabling the automated processing, analysis, and classification of vast amounts of textual data, NLP can help identify, filter, and correct misleading information while amplifying scientifically accurate content. This becomes especially important in low-resource language regions, such as Bengali (Bangla)-speaking communities, where access to accurate climate information is limited. Through NLP, improved information dissemination can reach the populations most affected by climate change, ensuring that they have the knowledge necessary to advocate for and implement more effective climate policies (Stede & Patz, 2021; Nicholls & Culpepper, 2020). NLP can also enhance the accessibility of global climate discussions, helping to break down language barriers that prevent these communities from engaging in critical debates and receiving the latest scientific developments (Nicholls & Culpepper, 2020).
@@
To overcome these issues, we present Dhoroni, a novel multi perspective Bangla (Bangali0 climate change and environmental news dataset. Dhoroni (pronounced "Dharaṇī") signifies "mother earth" in Bengali (Bangla), inspiring us to preserve our birthplace, Mother Earth, and protect the environment to prevent climate change. In addition to the dataset, we introduce a fmaily of baseline Bengali (Bangla) language models, named BanglaBERT-Dhoroni, fine-tuned on this dataset. This research gap creates a significant barrier to the effective processing and sharing of climate-related information from Bangladesh within the global scientific community. By underutilizing the potential of NLP for languages like Bengali (Bangla), the international conversation surrounding climate change risks becoming skewed, favoring voices from more developed, English-speaking nations. This further delays the recognition and action needed to address the environmental crises faced by nations like Bangladesh. Therefore, it is imperative to prioritize developing NLP technologies for low-resource languages to ensure a more equitable and inclusive approach to global climate solutions.
@@
    We introduce Dhoroni, a novel multi-perspective benchmark dataset, containing a collection of 2300 Bengali (Bangla) news articles from various sources and perspectives worldwide.
@@ 2 Related Works
The intersection of climate change and NLP has become increasingly important, with researchers exploring how NLP can contribute to understanding and addressing climate challenges. Stede & Patz (2021); Rolnick et al. (2022) emphasize the significance of integrating NLP with social science to navigate complex climate discussions, providing valuable insights for policy-making and activism. Mallick et al. (2024) highlight the role of NLP in understanding regional climate impacts, stressing the importance of localized adaptation and mitigation efforts. Furthermore, studies on climate awareness in NLP research emphasize the importance of fostering climate-conscious approaches within the NLP community, advocating for technology-driven solutions that contribute to global climate action (Hershcovich et al., 2022).
@@
Figure 2: Keyword used for data collection in Dhoroni dataset.
@@ 4 Model Design and Development
Model.  We have fine-tuned BanglaBERT (sagorsarker/bangla-bert-base) [^2] https://huggingface.co/sagorsarker/bangla-bert-base [^3] https://github.com/sagorbrur/bangla-bert, which is freely available in Hugging Face by transformers library (Wolf et al., 2020) with MIT Licence. It is a pretrained language model of Bengali language using mask language modeling described in BERT (Devlin et al., 2019).
@@
Task Location
Stance Detection ciol-research/BanglaBERT-Dhoroni-SD
News Authenticity Identification ciol-research/BanglaBERT-Dhoroni-NAD
Political Influence Detection ciol-research/BanglaBERT-Dhoroni-PAD
Scientific Data Usage Detection ciol-research/BanglaBERT-Dhoroni-SDU
Statistical Data Usage Detection ciol-research/BanglaBERT-Dhoroni-StDU
Data Sources Detection ciol-research/BanglaBERT-Dhoroni-DSD
Impacted Location or Countries Detection ciol-research/BanglaBERT-Dhoroni-ILD
Detecting Climate/Environmental Topics ciol-research/BanglaBERT-Dhoroni-CET
News Target Detection ciol-research/BanglaBERT-Dhoroni-NTD
Authority Involvement Detection ciol-research/BanglaBERT-Dhoroni-AID
@@ 5.1 Experimental Results and Discussion
No Task Accuracy ($\uparrow$) Precision ($\uparrow$) Recall ($\uparrow$) F1 ($\uparrow$)
@@
7 Impacted Location Detection 0.536 0.138 0.175 0.134
8 Detecting Climate/Env. Topics 0.052 0.015 0.076 0.017
@@
In Stance Detection, the accuracy of 0.635 suggests the model performs reasonably well overall, but the low precision (0.426) and recall (0.295) indicate it struggles to correctly identify stances, especially minority or less obvious ones. Stances, such as "support" or "oppose," can be subtle in language, and the model likely finds it difficult to differentiate between them, leading to more incorrect classifications.
Moving on to News Authenticity Identification, while the accuracy is impressively high at 0.911, precision (0.184) and recall (0.198) are extremely low. This is a classic sign of class imbalance: the model may be biased toward classifying most news as authentic, since it’s likely the majority class in the training data. As a result, it rarely identifies false news, which heavily skews its precision and recall. A more balanced dataset or targeted strategies to boost minority class detection could help improve these scores.
In Political Influence Detection, we see a better balance between precision (0.417) and recall (0.500), which is reflected in a decent F1 score of 0.455. This suggests the model is able to identify political influence with moderate success, though it still tends to miss some instances or incorrectly classify others. Political influence might vary significantly depending on context, which could explain why the model sometimes struggles to capture it consistently
Scientific Data Usage Detection performs relatively well, with an accuracy of 0.870 and a balanced F1 score of 0.465. The model seems to manage identifying scientific data usage, likely because this task involves more structured or identifiable cues (like specific terminology or phrasing). However, the model’s slight tendency to overestimate scientific usage (higher recall than precision) may lead it to incorrectly flag some cases.
In contrast, Statistical Data Usage Detection has lower accuracy (0.594), though its precision (0.491) and recall (0.497) are more balanced. This suggests the model can identify statistical data usage somewhat effectively, but the overall complexity of the task—where statistical data may appear in various forms—leads to less consistent results. The slight drop in F1 score (0.440) points to misclassification of some borderline cases.
Data Sources Detection shows moderate performance with an accuracy of 0.623, and relatively balanced precision and recall (both around 0.48). This task may be easier for the model to manage because references to data sources are often explicit, making it easier to pick up on relevant patterns. However, there may still be subtle variations in how sources are mentioned, leading to some misclassifications.
On the other hand, Impacted Location or Countries Detection shows poor performance. With a precision of 0.138 and an F1 score of 0.134, it’s clear the model struggles to consistently identify locations or countries affected by events. This might be due to the broad range of ways in which locations are mentioned in text, from indirect references to unclear geographical descriptions, making it harder for the model to extract this information reliably.
A particularly striking failure is seen in Detecting Climate/Environmental Topics, where the accuracy is just 0.052, and both precision and recall are near zero. This suggests the model is almost completely unable to identify climate-related content, likely due to either insufficient training data or highly ambiguous definitions of what constitutes a "climate" topic. It could also be that the model is not tuned to capture the wide variety of terms and context in which climate issues are discussed.
News Target Detection also shows significant challenges, with low precision (0.048) and a recall of only 0.125. Identifying the target of news content is inherently complex, often involving subtle contextual cues that the model seems unable to consistently capture. This might explain the low F1 score of 0.069, as the model struggles to pinpoint the correct targets amidst varying contexts.
Lastly, Authority Involvement Detection presents similarly poor results, with an accuracy of 0.183 and extremely low precision (0.030). Authority involvement may be indicated in a variety of ways—sometimes indirectly—which the model likely fails to grasp. The low recall (0.102) further confirms that the model is not identifying enough relevant instances, perhaps due to unclear definitions of authority involvement in the data.
@@ 6 Discussion
As awareness of environmental issues increases among governments and the public (Liu et al., 2023), our work with the Dhoroni dataset and the BanglaBERT-Dhoroni model family contributes significantly to this vital discourse. By providing a comprehensive resource that captures climate-related narratives in Bengali, we aim to facilitate informed discussions on sustainable environmental management. Our efforts not only empower environmental managers and stakeholers, but also engage a wider audience, enabling them to better understand and respond to the complexities of climate change (Ghimire & Panday, 2016) and its local implications.
@@ 6.1 Societal Impact
Societal impact of the Dhoroni dataset cannot be overstated. In Bangladesh, where climate change poses significant threats (Nishat & Mukherjee, 2013), understanding public discourse around these issues is crucial for driving social change (Bell, 1994). The dataset offers a platform for analyzing how climate-related information is communicated and perceived within the community. This understanding can empower citizens to advocate for more robust climate policies and hold leaders accountable for environmental actions.
By providing a rich repository of news articles, the dataset encourages grassroots movements focused on climate awareness and action. Local organizations can harness insights derived from the dataset to mobilize communities, fostering a collective sense of responsibility toward environmental conservation. This grassroots approach can create a ripple effect, leading to increased public engagement and advocacy for climate-related policies. Furthermore, the dataset’s emphasis on authenticity and data sources plays a pivotal role in combating misinformation. In an age where misinformation can easily spread through social media and news outlets (Allcott et al., 2019), having a reliable dataset to verify claims becomes essential. By highlighting credible sources and assessing the authenticity of information, the dataset can serve as a foundation for educating the public about the importance of fact-checking and responsible consumption of news (López-Marcos & Vicente-Fernández, 2021).
@@
Societal implications of the dataset also extend to gender and social equity issues. Climate change disproportionately affects marginalized communities, including women (Islam & Winkel, 2017), low-income households (Haque et al., 2014), and indigenous populations (Abate & Kronk, 2013). By analyzing media representations of these groups in the context of climate change, researchers can identify gaps in coverage and advocate for more inclusive narratives. This focus on equity is essential for ensuring that all voices are heard in the conversation about climate change and its impacts. Ultimately, the Dhoroni dataset has the potential to catalyze positive societal change by fostering informed discussions, encouraging activism, and promoting equity in climate discourse. By leveraging the insights generated from this dataset, individuals and organizations can work together to address climate challenges and create a more sustainable future.
@@ 6.2 Global Impact
Global impact of the Dhoroni dataset extends beyond the borders of Bangladesh, contributing to the broader conversation on climate change and environmental issues. As climate change is a global challenge (Bhat & Lone, 2023), understanding how different countries and cultures discuss these topics is vital. By providing a dataset focused on Bengali perspectives, the Dhoroni project adds an important dimension to the global discourse, highlighting the voices and experiences of a population that is often overlooked in international discussions.
The inclusion of Bengali-language content in global climate research is crucial for creating a more balanced narrative. Much of the existing climate-related research and discourse is dominated by English-language sources (Ghosh & Boykoff, 2019), which can skew the understanding of how climate change affects various regions and communities (Leal Filho et al., 2022). By amplifying Bengali voices, the dataset helps to ensure that the unique challenges and perspectives of Bangladeshis are recognized in global climate discussions. Additionally, the dataset can facilitate cross-cultural comparisons in climate reporting and public perception. Researchers and policymakers can analyze how climate change is framed in Bangladesh compared to other countries, revealing potential differences in priorities and narratives. This comparative analysis can inform international collaboration on climate initiatives, allowing for a more holistic understanding of the global climate landscape. On a global scale, BanglaBERT-Dhoroni family models can enhance understanding of how different regions are affected by news topics, such as environmental impacts or political issues. By detecting authenticity and target groups, they can help gauge the global relevance of news stories (Gurevitch & Kavoori, 1994) and ensure that critical issues, such as authority involvement and environmental topics, are accurately represented across different parts of the world.
@@
6.3 Dhoroni in Environmental Policy-making and Management
The Dhoroni dataset is a significant advancement in NLP for environmental management, especially for Bengali-speaking communities. By curating 2,300 climate-related news articles, it offers insights into how local media frames environmental issues, which is crucial for informed policymaking and public engagement (Sofiullahi et al., 2024). The fine-tuned BanglaBERT-Dhoroni models, leveraging transformer-based architectures, facilitate accurate detection of climate topics, authority involvement, and media perspectives. These insights help stakeholders address misinformation, promote awareness, and ensure marginalized voices are represented (Chen et al., 2019). Overall, the Dhoroni dataset and model family provide a powerful resource to enhance understanding of climate issues, empowering communities to proactively address climate challenges.